OpenStudy (anonymous):

INTEREST- REVISITED: Suzanne borrowed $820 from the bank for one year. if the annual interest rate was 15% what was the total amount she owed the bank at the end of the year??? show work PLZ!

OpenStudy (a_clan):

Principal amount (P) = $820 Interest rate (R) = 15% Time period (T) = 1 year Interest (I) = P*R*T = 820 * (15/100) * 1 =$123 Total amount (A) = P + I = $820+ $123 = $943 (Answer)